C&A Fast Food has four activities in serving a customer: greet customer, take order, process order, and deliver order. Each activity is staffed by one employee (for a total of four employees). The processing time for each activity is given as follows:

Activity

Processing time per customer

Greet customer

2 seconds

Take order

30 seconds

Process order

60 seconds

Deliver order

5 seconds

a.What is the average labor utilization?

b.Assume demand is unlimited. If one additional employee is added to the bottleneck activity, what will be the percentage change in process capacity?
Answer #1

Total capacity of Greet customer employee = 60 seconds / 2 = 30 customers / minute

Similarly, total capacity of Take order employee = 60/30 = 2 customers / minute

Process order = 60/60 = 1 customer / minute

Deliver order = 60 / 5 = 12 customers / minute

Here , bottleneck is processing order which determines the capacity of process i.e. 1 customer / minute.

Idle time =sun of ( bottleneck time - time for other processes)

Total idle time = (60-2) + (60-60) + (60-30) + (60-5) = 143 seconds / customer

Total task time = 2+30+60+5 = 97 seconds / customer

Average labor utilization = sum of task time /( sum of task time + idle time) = 97 / 240 = 40.416% or 40.42%

b) When we place another resource at bottleneck i.e. process order the effextive time to serve 1 customer = 60/2= 30 seconds

New process capacity = 60/ 30 = 2 customers / minute

Change in percentage process capacity = ( new process capacity - old capacity) / old capacity = (2-1)/ 1 = 100%